OpenStudy (anonymous):

solve each system by elimination Please help and solve y -2x-2y=6 10x+10y=-30

OpenStudy (anonymous):

You don't need to solve it. it is the same line

OpenStudy (anonymous):

huhh

OpenStudy (anonymous):

divide 10x+10y=-30 by 5 and get 2x+2y=-6 then multiply by negative 1 and get -2x-2y=6 which is the same as the first equation- they're the same eq so it's the same line. (just written differently)